The Mountain States Office of State Farm Insurance Company reports that approximately 73% of all automobile damage liability claims were made by people under 25 years of age. A random sample of twelve automobile insurance liability claims is under study.

(a) Make a histogram showing the probability that r = 0 to 12 claims are made by people under 25 years of age.

Maple Generated Plot Maple Generated Plot
Maple Generated Plot Maple Generated Plot


(b) Find the mean and standard deviation of this probability distribution. (Round your answers to two decimal places.)

μ =
σ =


For samples of size 12, what is the expected number of claims made by people under 25 years of age? (Round your answer to the nearest whole number.)
Answer #1

This will be a Binomial distribution with parameters:

n = 12, p = 0.73

Hence,

a) Since the probability is greater than 0.5, this distribution will be left skewed. Hence,

Graph B is correct.

b) Mean = np = 12*0.73 = 8.76

Standard deviation = pq = V12 * 0.73 * 0.27 = 1.54

Expected number of claims made = 9